# FINITE-DIMENSIONAL GAUSSIAN APPROXIMATION FOR DEEP NEURAL NETWORKS: UNIVERSALITY IN RANDOM WEIGHTS

### v1

@@ 2.2 Error for Gaussian weights to Gaussian
$\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(F_{1}^{(L-1)}(x))^{2p}\bigr{]}$ $=\sum_{j_{1},\ldots,j_{2p}=1}^{n_{L-2}}\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\prod_{\ell=1}^{2p}W_{1,j_{\ell}}^{(L-2)}\biggr{]}\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\prod_{\ell=1}^{2p}\sigma(F^{(L-2)}_{j_{\ell}}(x))\biggr{]}$
$=\sum_{(j_{1},\ldots,j_{2p})\in A_{n_{L-2}}^{(2p)}}\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\prod_{\ell=1}^{2p}W_{1,j_{\ell}}^{(L-2)}\biggr{]}\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\prod_{\ell=1}^{2p}\sigma(F^{(L-2)}_{j_{\ell}}(x))\biggr{]}$
$\leqslant\bigl{\lvert}A_{n_{L-2}}^{(2p)}\bigr{\rvert}\frac{c_{2p}^{(L-2)}}{n_{L-2}^{p}}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(F_{1}^{(L-2)}(x))\bigr{)}^{2p}\bigr{]}$
$\leqslant c\cdot c_{2p}^{(L-2)}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(F_{1}^{(L-2)}(x))\bigr{)}^{2p}\bigr{]},$
@@
$\begin{split}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(F_{1}^{(L-1)}(x))\bigr{)}^{2p}\bigr{]}&\leqslant\bigl{(}\mathrm{Lip}_{\sigma}\cdot c\cdot c_{2p}^{(L-2)}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(F_{1}^{(L-2)}(x))\bigr{)}^{2p}\bigr{]}+\lvert\sigma(0)\rvert\bigr{)}^{2p}\\ &\leqslant(\mathrm{Lip}_{\sigma}+\lvert\sigma(0)\rvert)^{2p}\bigl{\{}c\cdot c_{2p}^{(L-2)}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(F_{1}^{(L-2)}(x))\bigr{)}^{2p}\bigr{]}\vee 1\bigr{\}}.\end{split}$ (2.16)
@@
$\begin{split}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(F_{1}^{(1)}(x))^{2p}\bigr{]}&=\sum_{j_{1},\ldots,j_{2p}=1}^{n_{L-2}}\mathop{{}\mathbbm{E}}\mathopen{}\biggl{[}\prod_{\ell=1}^{2p}W_{1,j_{\ell}}^{(0)}\biggr{]}\prod_{\ell=1}^{2p}x_{j_{\ell}}\leqslant\frac{c_{2p}^{(0)}}{n_{0}^{p}}\sum_{j_{1},\ldots,j_{2p}=1}^{n_{L-2}}\ \prod_{\ell=1}^{2p}\lvert x_{j_{\ell}}\rvert=\frac{c_{2p}^{(0)}}{n_{0}^{p}}\|x\|_{1}^{2p}.\end{split}$ (2.17)
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(G_{1}^{(L-1)}(x))\bigr{)}^{2p}\bigr{]}^{1/(2p)}$ $\leqslant\mathrm{Lip}_{\sigma}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(G_{1}^{(L-1)}(x))^{2p}\bigr{]}^{1/(2p)}+\lvert\sigma(0)\rvert.$
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(G_{1}^{(L-1)}(x))^{2p}\bigr{]}$ $=(c_{w}^{(L-2)})^{p}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(G_{1}^{(L-1)}(x))^{2}\bigr{]}^{p}2^{p}\biggl{(}\frac{\Gamma(p+1/2)}{\Gamma(1/2)}\biggr{)}$
$\leqslant c\cdot(c_{w}^{(L-2)})^{p}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(G_{1}^{(L-2)}(x))^{p}\bigr{]}.$
@@
$\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\bigl{(}\sigma(G_{1}^{(L-1)}(x))\bigr{)}^{2p}\bigr{]}$ $\leqslant(\mathrm{Lip}_{\sigma}+\lvert\sigma(0)\rvert)^{2p}\bigl{\{}\bigl{(}c\cdot c_{2p}^{(L-2)}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(G_{1}^{(L-2)}(x))^{2}\bigr{]}^{p}\bigr{)}\vee 1\bigr{\}}.$
@@
Retaining the notation above, if there is an integer $p>2$ and constants $c_{2p}^{(\ell)}\geqslant 1$ for $\ell=0,\ldots,L-2$ such that $\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(W_{ij}^{(\ell)})^{2p}\bigr{]}\leqslant c_{2p}^{(\ell)}/n_{\ell}^{p}$, then for any $x_{a},x_{b}\in\mathcal{M}$, $\chi_{a,b}=(x_{a},x_{b})$, and $B^{(L-1)}_{2p}$ defined in Lemma 2.7, we have
$\begin{split}\Bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(G_{1}^{(L-1)}(x_{a}))\sigma(G_{1}^{(L-1)}&(x_{b}))\bigr{]}-\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(F_{1}^{(L-1)}(x_{a}))\sigma(F_{1}^{(L-1)}(x_{b}))\bigr{]}\Bigr{\rvert}\\ &\leqslant C_{2p}^{(L-1)}(x_{a},x_{b})\mathsf{d}_{1}\bigl{(}F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr{)}^{(2p-2)/(2p-1)},\end{split}$ (2.18)
@@
For the first bound (2.18), we apply Lemma 2.6 with $d=2$, $q=2p-2$, $X_{1}=\sigma(F^{(L-1)}_{1}(x_{a}))$, $X_{2}=\sigma(F_{1}^{(L-1)}(x_{b}))$, $Z_{1}=\sigma(G^{(L-1)}_{1}(x_{a}))$, $Z_{2}=\sigma(G_{1}^{(L-1)}(x_{b}))$, and, using Lemma 2.7, with
@@
$\begin{split}&\leqslant\Bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(F_{1}^{(L-1)}(x_{a}))\sigma(F_{1}^{(L-1)}(x_{b}))\sigma(F_{2}^{(L-1)}(x_{a}))\sigma(F_{2}^{(L-1)}(x_{b}))\bigr{]}\\ &\qquad\qquad-\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(G_{1}^{(L-1)}(x_{a}))\sigma(G_{1}^{(L-1)}(x_{b}))\sigma(G_{2}^{(L-1)}(x_{a}))\sigma(G_{2}^{(L-1)}(x_{b}))\bigr{]}\Bigr{\rvert}\end{split}$ (2.23)
$\begin{split}&\quad+\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(G_{1}^{(L-1)}(x_{a}))\sigma(G_{1}^{(L-1)}(x_{b}))\bigr{]}-\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(F_{1}^{(L-1)}(x_{a}))\sigma(F_{1}^{(L-1)}(x_{b}))\bigr{]}\bigr{\rvert}\\ &\qquad\qquad\times\bigl{(}\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(G_{1}^{(L-1)}(x_{a}))\sigma(G_{1}^{(L-1)}(x_{b}))\bigr{]}\bigr{\rvert}+\bigl{\lvert}\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}\sigma(F_{1}^{(L-1)}(x_{a}))\sigma(F_{1}^{(L-1)}(x_{b}))\bigr{]}\bigr{\rvert}\bigr{)}.\end{split}$ (2.24)
We can bound (LABEL:eq:4tfg) using Lemma 2.6 with $d=4$, $q=2p-4$, $X_{1},X_{2},Z_{1},Z_{2}$ as above, and $X_{3},X_{4},Z_{3},Z_{4}$ defined as $X_{1},X_{2},Z_{1},Z_{2}$, but with second components instead of first, and using Lemma 2.7, with
@@
Retaining the notation above, assume there is an integer $p>2$ and constants $c_{2p}^{(\ell)}\geqslant 1$ for $\ell=0,\ldots,L-2$ such that $\mathop{{}\mathbbm{E}}\mathopen{}\bigl{[}(W_{ij}^{(\ell)})^{2p}\bigr{]}\leqslant c_{2p}^{(\ell)}/n_{\ell}^{p}$, and for any $a,b\in\{1,\ldots,s\}$ write $\chi_{a,b}=(x_{a},x_{b})$. Then
$\mathsf{d}_{3}(\widetilde{F}_{[d]}^{(L)},G_{[d]}^{(L)})$
$\leqslant\frac{c_{2}^{(L-1)}\,d}{2}\sum_{a,b=1}^{s}\biggl{\{}C_{2p}^{(L-1)}(x_{a},x_{b})\mathsf{d}_{1}\bigl{(}F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr{)}^{(2p-2)/(2p-1)}$
$\hskip 99.58464pt+\frac{\bigl{(}B_{2p}^{(L-1)}(x_{a})B_{2p}^{(L-1)}(x_{b})\bigr{)}^{1/(2p)}}{\sqrt{n_{L-1}}}$
$\hskip 99.58464pt+2C_{2p}^{(L-1)}(x_{a},x_{b})\sqrt{B_{2}(x_{a})B_{2}(x_{b})}\mathsf{d}_{1}\bigl{(}F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr{)}^{(p-1)/(2p-1)}$
$\hskip 99.58464pt+D_{2p}^{(L-1)}(x_{a},x_{b})\mathsf{d}_{1}\bigl{(}F^{(L-1)}_{[2]}(\chi_{a,b}),G^{(L-1)}_{[2]}(\chi_{a,b})\bigr{)}^{(p-2)/(2p-1)}\biggr{\}},$
@@
$\mathsf{d}_{3}\bigl{(}F^{(L)}(\chi),G^{(L)}(\chi)\bigr{)}$ (2.27)
@@
$\hskip 99.58464pt+\frac{\bigl{(}B_{2p}^{(L-1)}(x_{a})B_{2p}^{(L-1)}(x_{b})\bigr{)}^{1/(2p)}}{\sqrt{n_{L-1}}}$
$\hskip 99.58464pt+2C_{2p}^{(L-1)}(x_{a},x_{b})\sqrt{B_{2}(x_{a})B_{2}(x_{b})}\mathsf{d}_{1}\bigl{(}F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr{)}^{(p-1)/(2p-1)}$
$\hskip 99.58464pt+D_{2p}^{(L-1)}(x_{a},x_{b})\mathsf{d}_{1}\bigl{(}F^{(L-1)}_{[2]}(\chi_{a,b}),G^{(L-1)}_{[2]}(\chi_{a,b})\bigr{)}^{(p-2)/(2p-1)}\biggr{\}}$
$\qquad+\frac{n_{L}}{\sqrt{n_{L-1}}}\biggl{(}\frac{2(c_{2}^{(L-1)})^{3/2}+c_{3}^{(L-1)}}{6}\biggr{)}\sum_{a=1}^{s}\mathop{{}\mathbbm{E}}\mathopen{}\Bigl{[}\bigl{\lvert}\sigma(F_{1}^{(L-1)}(x_{a})\bigr{\rvert}^{3}\Bigr{]}$
$\hskip 99.58464pt+2C_{2p}^{(L-1)}(x_{a},x_{b})\sqrt{B_{2}(x_{a})B_{2}(x_{b})}\mathsf{d}_{1}\bigl{(}F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr{)}^{(p-1)/(2p-1)}$
$\hskip 99.58464pt+D_{2p}^{(L-1)}(x_{a},x_{b})\mathsf{d}_{1}\bigl{(}F^{(L-1)}_{[2]}(\chi_{a,b}),G^{(L-1)}_{[2]}(\chi_{a,b})\bigr{)}^{(p-2)/(2p-1)}\biggr{\}}$
$\qquad+\frac{n_{L}}{\sqrt{n_{L-1}}}\biggl{\{}\biggl{(}\frac{2(c_{2}^{(L-1)})^{3/2}+c_{3}^{(L-1)}}{6}\biggr{)}\sum_{a=1}^{s}B_{2p}^{(L-1)}(x_{a})^{3/(2p)}+\frac{c_{2}^{(L-1)}}{2}\biggl{(}\sum_{a=1}^{s}B_{2p}^{(L-1)}(x_{a})\biggr{)}^{2}\biggr{\}}.$
@@
$\mathsf{d}_{3}\bigl{(}F^{(2)}_{[d]}(\chi_{a,b}),G_{[d]}^{(2)}(\chi_{a,b})\bigr{)}$ $\leqslant\frac{d}{\sqrt{n_{1}}}\biggl{(}\frac{2(c_{2}^{(1)})^{3/2}+c_{3}^{(1)}}{6}\biggr{)}\sum_{c\in\{a,b\}}\mathop{{}\mathbbm{E}}\mathopen{}\Bigl{[}\bigl{\lvert}\sigma(F_{1}^{(1)}(x_{c})\bigr{\rvert}^{3}\Bigr{]}$
@@
$\begin{split}\mathsf{d}_{3}&\bigl{(}F^{(\ell+1)}_{[d]}(\chi_{a,b}),G_{[d]}^{(\ell+1)}(\chi_{a,b})\bigr{)}\\ &\leqslant\frac{c_{2}^{(\ell)}\,d}{2}\biggl{\{}C_{2p}^{(\ell)}(x_{a},x_{b})\mathsf{d}_{1}\bigl{(}F^{(\ell)}_{[1]}(\chi_{a,b}),G^{(\ell)}_{[1]}(\chi_{a,b})\bigr{)}^{(2p-2)/(2p-1)}\\ &\hskip 113.81102pt+2C_{2p}^{(\ell)}(x_{a},x_{b})\sqrt{B_{2}^{(\ell)}(x_{a})B_{2}^{(\ell)}(x_{b})}\mathsf{d}_{1}\bigl{(}F^{(\ell)}_{[1]}(\chi_{a,b}),G^{(\ell)}_{[1]}(\chi_{a,b})\bigr{)}^{(p-1)/(2p-1)}\\ &\hskip 113.81102pt+D_{2p}^{(\ell)}(x_{a},x_{b})\mathsf{d}_{1}\bigl{(}F^{(\ell)}_{[2]}(\chi_{a,b}),G^{(\ell)}_{[2]}(\chi_{a,b})\bigr{)}^{(p-2)/(2p-1)}\biggr{\}}\\ &\qquad\qquad+\frac{d}{\sqrt{n_{\ell}}}\biggl{(}\frac{2(c_{2}^{(\ell)})^{3/2}+6c_{2}^{(\ell)}+c_{3}^{(\ell)}}{6}\biggr{)}\sum_{c\in\{a,b\}}B_{2p}^{(\ell)}(x_{c})^{3/(2p)}\\ &\leqslant c_{2}^{(\ell)}\biggl{\{}C_{2p}^{(\ell)}(x_{a},x_{b})\bigl{(}1+2\sqrt{B_{2}^{(\ell)}(x_{a})B_{2}^{(\ell)}(x_{b})}\bigr{)}+D_{2p}^{(\ell)}(x_{a},x_{b})\biggr{\}}\\ &\hskip 71.13188pt\times\Bigl{(}\mathtt{C}_{\ell}\sum_{m=1}^{\ell-1}n_{m}^{-\frac{1}{6}\left(\frac{p-2}{3(2p-1)}\right)^{\ell-1-m}}\Bigr{)}^{(p-2)/(2p-1)}\\ &\qquad\qquad+\frac{2}{\sqrt{n_{\ell}}}\biggl{(}\frac{2(c_{2}^{(\ell)})^{3/2}+6c_{2}^{(\ell)}+c_{3}^{(\ell)}}{6}\biggr{)}\sum_{c\in\{a,b\}}B_{2p}^{(\ell)}(x_{c})^{3/(2p)},\end{split}$ (2.29)

### v2

@@ 2.2 Error for Gaussian weights to Gaussian
$\mathbbm{E}\bigl[(F_{1}^{(L-1)}(x))^{2p}\bigr]$ $=\sum_{j_{1},\ldots,j_{2p}=1}^{n_{L-2}}\mathbbm{E}\biggl[\prod_{\ell=1}^{2p}W_{1,j_{\ell}}^{(L-2)}\biggr]\mathbbm{E}\biggl[\prod_{\ell=1}^{2p}\sigma(F^{(L-2)}_{j_{\ell}}(x))\biggr]$
$=\sum_{(j_{1},\ldots,j_{2p})\in A_{n_{L-2}}^{(2p)}}\mathbbm{E}\biggl[\prod_{\ell=1}^{2p}W_{1,j_{\ell}}^{(L-2)}\biggr]\mathbbm{E}\biggl[\prod_{\ell=1}^{2p}\sigma(F^{(L-2)}_{j_{\ell}}(x))\biggr]$
$\leqslant\bigl\lvert A_{n_{L-2}}^{(2p)}\bigr\rvert\frac{c_{2p}^{(L-2)}}{n_{L-2}^{p}}\mathbbm{E}\bigl[\bigl(\sigma(F_{1}^{(L-2)}(x))\bigr)^{2p}\bigr]$
$\leqslant c\cdot c_{2p}^{(L-2)}\mathbbm{E}\bigl[\bigl(\sigma(F_{1}^{(L-2)}(x))\bigr)^{2p}\bigr],$
@@
$\begin{split}\mathbbm{E}\bigl[\bigl(\sigma(F_{1}^{(L-1)}(x))\bigr)^{2p}\bigr]&\leqslant\Bigl(\mathrm{Lip}_{\sigma}\bigl(c\cdot c_{2p}^{(L-2)}\mathbbm{E}\bigl[\bigl(\sigma(F_{1}^{(L-2)}(x))\bigr)^{2p}\bigr]\bigr)^{1/(2p)}+\lvert\sigma(0)\rvert\Bigr)^{2p}\\ &\leqslant\bigl(\mathrm{Lip}_{\sigma}+\lvert\sigma(0)\rvert\bigr)^{2p}\Bigl\{\bigl(c\cdot c_{2p}^{(L-2)}\mathbbm{E}\bigl[\bigl(\sigma(F_{1}^{(L-2)}(x))\bigr)^{2p}\bigr]\bigr)\vee 1\Bigr\}.\end{split}$ (2.17)
@@
$\begin{split}\mathbbm{E}\bigl[(F_{1}^{(1)}(x))^{2p}\bigr]&=\sum_{j_{1},\ldots,j_{2p}=1}^{n_{L-2}}\mathbbm{E}\biggl[\prod_{\ell=1}^{2p}W_{1,j_{\ell}}^{(0)}\biggr]\prod_{\ell=1}^{2p}x_{j_{\ell}}\leqslant\frac{c_{2p}^{(0)}}{n_{0}^{p}}\sum_{j_{1},\ldots,j_{2p}=1}^{n_{L-2}}\ \prod_{\ell=1}^{2p}\lvert x_{j_{\ell}}\rvert=\frac{c_{2p}^{(0)}}{n_{0}^{p}}\|x\|_{1}^{2p}.\end{split}$ (2.18)
@@
$\mathbbm{E}\bigl[\bigl(\sigma(G_{1}^{(L-1)}(x))\bigr)^{2p}\bigr]^{1/(2p)}$ $\leqslant\mathrm{Lip}_{\sigma}\mathbbm{E}\bigl[(G_{1}^{(L-1)}(x))^{2p}\bigr]^{1/(2p)}+\lvert\sigma(0)\rvert$.
@@
$\mathbbm{E}\bigl[(G_{1}^{(L-1)}(x))^{2p}\bigr]$ $=(c_{w}^{(L-2)})^{p}\mathbbm{E}\bigl[(G_{1}^{(L-1)}(x))^{2}\bigr]^{p}2^{p}\biggl(\frac{\Gamma(p+1/2)}{\Gamma(1/2)}\biggr)$
$\leqslant c\cdot(c_{w}^{(L-2)})^{p}\mathbbm{E}\bigl[\sigma(G_{1}^{(L-2)}(x))^{p}\bigr].$
@@
$\mathbbm{E}\bigl[\bigl(\sigma(G_{1}^{(L-1)}(x))\bigr)^{2p}\bigr]$ $\leqslant(\mathrm{Lip}_{\sigma}+\lvert\sigma(0)\rvert)^{2p}\bigl\{\bigl(c\cdot c_{2p}^{(L-2)}\mathbbm{E}\bigl[\sigma(G_{1}^{(L-2)}(x))^{2}\bigr]^{p}\bigr)\vee 1\bigr\}.$
@@
Retaining the notation above, if there is an integer $p>2$ and constants $c_{2p}^{(\ell)}\geqslant 1$ for $\ell=0,\ldots,L-2$ such that $\mathbbm{E}\bigl[(W_{ij}^{(\ell)})^{2p}\bigr]\leqslant c_{2p}^{(\ell)}/n_{\ell}^{p}$, then for any $x_{a},x_{b}\in\mathcal{M}$, $\chi_{a,b}=(x_{a},x_{b})$, and $B^{(L-1)}_{2p}$ defined in Lemma 2.7, we have
$\begin{split}\Bigl\lvert\mathbbm{E}\bigl[\sigma(G_{1}^{(L-1)}(x_{a}))\sigma(G_{1}^{(L-1)}&(x_{b}))\bigr]-\mathbbm{E}\bigl[\sigma(F_{1}^{(L-1)}(x_{a}))\sigma(F_{1}^{(L-1)}(x_{b}))\bigr]\Bigr\rvert\\ &\leqslant C_{2p}^{(L-1)}(x_{a},x_{b})\mathsf{d}_{1}\bigl(F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr)^{(2p-2)/(2p-1)},\end{split}$ (2.19)
@@
For the first bound (2.19), we apply Lemma 2.6 with $d=2$, q=2p2q=2p-2, $X_{1}=\sigma(F^{(L-1)}_{1}(x_{a}))$, $X_{2}=\sigma(F_{1}^{(L-1)}(x_{b}))$, $Z_{1}=\sigma(G^{(L-1)}_{1}(x_{a}))$, $Z_{2}=\sigma(G_{1}^{(L-1)}(x_{b}))$, and, using Lemma 2.7, with
@@
$\begin{split}&\leqslant\Bigl\lvert\mathbbm{E}\bigl[\sigma(F_{1}^{(L-1)}(x_{a}))\sigma(F_{1}^{(L-1)}(x_{b}))\sigma(F_{2}^{(L-1)}(x_{a}))\sigma(F_{2}^{(L-1)}(x_{b}))\bigr]\\ &\qquad\qquad-\mathbbm{E}\bigl[\sigma(G_{1}^{(L-1)}(x_{a}))\sigma(G_{1}^{(L-1)}(x_{b}))\sigma(G_{2}^{(L-1)}(x_{a}))\sigma(G_{2}^{(L-1)}(x_{b}))\bigr]\Bigr\rvert\end{split}$ (2.24)
[…]
We can bound (LABEL:eq:4tfg) using Lemma 2.6 with $d=4$, q=2p4q=2p-4, $X_{1},X_{2},Z_{1},Z_{2}$ as above, and $X_{3},X_{4},Z_{3},Z_{4}$ defined as $X_{1},X_{2},Z_{1},Z_{2}$, but with second components instead of first, and using Lemma 2.7, with
@@
Retaining the notation above, assume there is an integer $p>2$ and constants $c_{2p}^{(\ell)}\geqslant 1$ for $\ell=0,\ldots,L-2$ such that $\mathbbm{E}\bigl[(W_{ij}^{(\ell)})^{2p}\bigr]\leqslant c_{2p}^{(\ell)}/n_{\ell}^{p}$, and for any $a,b\in\{1,\ldots,s\}$ write $\chi_{a,b}=(x_{a},x_{b})$. Then
$\mathsf{d}_{3}\bigl(\widetilde{F}_{[d]}^{(L)}(\chi),G_{[d]}^{(L)}(\chi)\bigr)$ $\leqslant\frac{c_{2}^{(L-1)}\,d}{2}\sum_{a,b=1}^{s}\biggl\{C_{2p}^{(L-1)}(x_{a},x_{b})\mathsf{d}_{1}\bigl(F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr)^{\frac{(2p-2)}{(2p-1)}}$
$\hskip 5.69054pt+\frac{\bigl(B_{2p}^{(L-1)}(x_{a})B_{2p}^{(L-1)}(x_{b})\bigr)^{1/(2p)}}{\sqrt{n_{L-1}}}$
$\hskip 5.69054pt+\sqrt{2C_{2p}^{(L-1)}(x_{a},x_{b})\sqrt{B_{2}(x_{a})B_{2}(x_{b})}}\mathsf{d}_{1}\bigl(F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr)^{\frac{(p-1)}{(2p-1)}}$
$\hskip 5.69054pt+\sqrt{D_{2p}^{(L-1)}(x_{a},x_{b})}\mathsf{d}_{1}\bigl(F^{(L-1)}_{[2]}(\chi_{a,b}),G^{(L-1)}_{[2]}(\chi_{a,b})\bigr)^{(p-2)/(2p-1)}\biggr\},$
@@
$\hskip 14.22636pt\mathsf{d}_{3}\bigl(F^{(L)}(\chi),G^{(L)}(\chi)\bigr)$
$\leqslant\frac{c_{2}^{(L-1)}\,n_{L}}{2}\sum_{a,b=1}^{s}\biggl\{C_{2p}^{(L-1)}(x_{a},x_{b})\mathsf{d}_{1}\bigl(F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr)^{\frac{(2p-2)}{(2p-1)}}$
$\hskip 5.69054pt+\frac{\bigl(B_{2p}^{(L-1)}(x_{a})B_{2p}^{(L-1)}(x_{b})\bigr)^{1/(2p)}}{\sqrt{n_{L-1}}}$
$\hskip 5.69054pt+\sqrt{2C_{2p}^{(L-1)}(x_{a},x_{b})\sqrt{B_{2}(x_{a})B_{2}(x_{b})}}\mathsf{d}_{1}\bigl(F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr)^{(p-1)/(2p-1)}$
$\hskip 5.69054pt+\sqrt{D_{2p}^{(L-1)}(x_{a},x_{b})}\mathsf{d}_{1}\bigl(F^{(L-1)}_{[2]}(\chi_{a,b}),G^{(L-1)}_{[2]}(\chi_{a,b})\bigr)^{(p-2)/(2p-1)}\biggr\}$
$\qquad+\frac{n_{L}\,s^{2}}{\sqrt{n_{L-1}}}\biggl(\frac{2(c_{2}^{(L-1)})^{3/2}+c_{3}^{(L-1)}}{6}\biggr)\sum_{a=1}^{s}\mathbbm{E}\Bigl[\bigl\lvert\sigma(F_{1}^{(L-1)}(x_{a})\bigr\rvert^{3}\Bigr]$
@@
$\hskip 5.69054pt+\sqrt{2C_{2p}^{(L-1)}(x_{a},x_{b})\sqrt{B_{2}(x_{a})B_{2}(x_{b})}}\mathsf{d}_{1}\bigl(F^{(L-1)}_{[1]}(\chi_{a,b}),G^{(L-1)}_{[1]}(\chi_{a,b})\bigr)^{(p-1)/(2p-1)}$
$\hskip 5.69054pt+\sqrt{D_{2p}^{(L-1)}(x_{a},x_{b})}\mathsf{d}_{1}\bigl(F^{(L-1)}_{[2]}(\chi_{a,b}),G^{(L-1)}_{[2]}(\chi_{a,b})\bigr)^{(p-2)/(2p-1)}\biggr\}$
$\qquad+\frac{n_{L}\,s^{2}}{\sqrt{n_{L-1}}}\biggl\{\biggl(\frac{2(c_{2}^{(L-1)})^{3/2}+c_{3}^{(L-1)}}{6}\biggr)\sum_{a=1}^{s}B_{2p}^{(L-1)}(x_{a})^{3/(2p)}+\frac{c_{2}^{(L-1)}}{2}\biggl(\sum_{a=1}^{s}B_{2p}^{(L-1)}(x_{a})\biggr)^{2}\biggr\}.$
@@
$\mathsf{d}_{3}\bigl(F^{(2)}_{[d]}(\chi_{a,b}),G_{[d]}^{(2)}(\chi_{a,b})\bigr)$ $\leqslant\frac{d}{\sqrt{n_{1}}}\biggl(\frac{2(c_{2}^{(1)})^{3/2}+c_{3}^{(1)}}{6}\biggr)\sum_{c\in\{a,b\}}\mathbbm{E}\Bigl[\bigl\lvert\sigma(F_{1}^{(1)}(x_{c})\bigr\rvert^{3}\Bigr]$
@@
$\begin{split}\mathsf{d}_{3}&\bigl(F^{(\ell+1)}_{[d]}(\chi_{a,b}),G_{[d]}^{(\ell+1)}(\chi_{a,b})\bigr)\\ &\leqslant\frac{c_{2}^{(\ell)}\,d}{2}\biggl\{C_{2p}^{(\ell)}(x_{a},x_{b})\mathsf{d}_{1}\bigl(F^{(\ell)}_{[1]}(\chi_{a,b}),G^{(\ell)}_{[1]}(\chi_{a,b})\bigr)^{(2p-2)/(2p-1)}\\ &\hskip 56.9055pt+\sqrt{2C_{2p}^{(\ell)}(x_{a},x_{b})\sqrt{B_{2}^{(\ell)}(x_{a})B_{2}^{(\ell)}(x_{b})}}\mathsf{d}_{1}\bigl(F^{(\ell)}_{[1]}(\chi_{a,b}),G^{(\ell)}_{[1]}(\chi_{a,b})\bigr)^{(p-1)/(2p-1)}\\ &\hskip 56.9055pt+\sqrt{D_{2p}^{(\ell)}(x_{a},x_{b})}\mathsf{d}_{1}\bigl(F^{(\ell)}_{[2]}(\chi_{a,b}),G^{(\ell)}_{[2]}(\chi_{a,b})\bigr)^{(p-2)/(2p-1)}\biggr\}\\ &\qquad\qquad+\frac{d}{\sqrt{n_{\ell}}}\biggl(\frac{2(c_{2}^{(\ell)})^{3/2}+6c_{2}^{(\ell)}+c_{3}^{(\ell)}}{6}\biggr)\sum_{c\in\{a,b\}}B_{2p}^{(\ell)}(x_{c})^{3/(2p)}\\ &\leqslant c_{2}^{(\ell)}\biggl\{C_{2p}^{(\ell)}(x_{a},x_{b})+\sqrt{2C_{2p}^{(\ell)}(x_{a},x_{b})\sqrt{B_{2}^{(\ell)}(x_{a})B_{2}^{(\ell)}(x_{b})}}+\sqrt{D_{2p}^{(\ell)}(x_{a},x_{b})}\biggr\}\\ &\hskip 71.13188pt\times\Bigl(\mathtt{C}_{\ell}\sum_{m=1}^{\ell-1}n_{m}^{-\frac{1}{6}\left(\frac{p-2}{3(2p-1)}\right)^{\ell-1-m}}\Bigr)^{(p-2)/(2p-1)}\\ &\qquad\qquad+\frac{2}{\sqrt{n_{\ell}}}\biggl(\frac{2(c_{2}^{(\ell)})^{3/2}+6c_{2}^{(\ell)}+c_{3}^{(\ell)}}{6}\biggr)\sum_{c\in\{a,b\}}B_{2p}^{(\ell)}(x_{c})^{3/(2p)},\end{split}$ (2.29)